OpenStudy (anonymous):

-3(a-2)+1=28, simplify and solve for ''a'' please:]

OpenStudy (anonymous):

\[\Large \begin{array}{l} - 3(a - 2) + 1 = 28\\ - 3a + 6 + 1 = 28\\ - 3a = 21\\ a = - 7 \end{array}\]

OpenStudy (anonymous):

Correct, a = -7.